OpenStudy (anonymous):

Use the binomial theorem to compute the square root of 26. (Hint = 26^1/2)

OpenStudy (anonymous):

A number raised to the 1/2 power is the same as taking the square root of the number.

OpenStudy (anonymous):

we know 5^2 = 25 which is close to 26 so we can write 26 = 25 + 1 then (26)^1/2 =(25 + 1)^1/2 = 25^1/2 (1 + 1/25)^1/2 = 5(1+1/25)^1/2 you can continue expanding to be more precise.

OpenStudy (anonymous):

@ktm thank you